# Deblur4DGS: 4D Gaussian Splatting from Blurry Monocular Video

### v2

@@ Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we take 3DGS as the scene representation manner, and propose Deblur4DGS to reconstruct a high-quality 4D model from blurry monocular video. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce the exposure regularization term, multi-frame, and multi-resolution consistency regularization term to avoid trivial solutions. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments in both synthetic and real-world data on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes will be publicly available.
@@ Introduction
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields (Zhu et al. 2024; Yang et al. 2024; Wu et al. 2024; Yan, Li, and Lee 2023) and explicit deformation (Duan et al. 2024b; Chu, Ke, and Fragkiadaki 2024; Katsumata, Vo, and Nakayama 2024; Lin et al. 2024b; Li et al. 2024; Wang et al. 2024) are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Zhu et al. 2024), tracks (Seidenschwarz et al. 2024; Lei et al. 2024), and generative models (Wu et al. 2025; Chu, Ke, and Fragkiadaki 2024; Zeng et al. 2025) for better 4D reconstruction.
Unfortunately, motion blur often arises due to camera shake and object movement. When reconstructing the 4D scene from the blurry video, the above methods usually render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based (Lee et al. 2023a; Wang et al. 2023; Lee et al. 2023b) and 3DGS-based (Zhao, Wang, and Liu 2024; Chen and Liu 2024; Oh et al. 2024) methods have suggested jointly optimizing 3D representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames.
In this work, we take 3DGS (Kerbl et al. 2023) as the scene representation manner to explore the problem, driven by two main motivations. First, its successful application in 4D reconstruction make this method highly promising. Second, the explicit 3D motion modeling presents an opportunity to simplify the complex continuous dynamic representations estimation within the exposure time into exposure time estimation, avoiding complex extra motion modeling in DyBluRF (Sun et al. 2024a; Bui et al. 2023). Once the exposure time is estimated, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps. We note that the concurrent work BARD-GS (Lu et al. 2025) adopts a similar strategy, but they perform unsatisfactorily due to the under-constrained optimization, especially for large object motion.
@@
Deep learning-based image (Ren et al. 2023; Li et al. 2023; Kong et al. 2023; Wang et al. 2022; Zhang et al. 2022; Lin et al. 2024a; Zhang et al. 2024) and video (Zhong et al. 2023b; Liang et al. 2024; Pan et al. 2023; Zhang, Xie, and Yao 2024; Zhong et al. 2023a, 2020; Chan et al. 2022) deblurring methods have been widely explored. Compared to image deblurring methods, video ones leverage temporal clues between consecutive frames for more effective restoration. DBLRNet (Zhang et al. 2018a) employs spatial-temporal 3D convolution for spatial and temporal features aggregation. DSTNet (Pan et al. 2023) develops a deep discriminative spatial and temporal network. BasicVSR++ (Chan et al. 2022) improves feature fusion with second-order feature propagation and flow-guided alignment. BSSTNet (Zhang, Xie, and Yao 2024) introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
@@
To reconstruct 3D models, NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results in both image quality and rendering speed. To reconstruct 4D models, most works (Somraj et al. 2024; Duan et al. 2024a; Lu et al. 2024; Luiten et al. 2023; Kratimenos, Lei, and Daniilidis 2023; Huang et al. 2024; Lin et al. 2024b; Li et al. 2024; Shaw et al. ; Sun et al. 2024b; Wu et al. 2024; Yang et al. 2024; Bae et al. 2024; Mihajlovic et al. 2024; Duan et al. 2024b) incorporate implicit neural fields and explicit deformation for motion representation. For example, DeformableGS (Yang et al. 2024) achieves it with multi-layer perceptrons. 4DGasusians (Bae et al. 2024) improves it with a multi-resolution HexPlane (Cao and Johnson 2023). E-D3DGS (Wu et al. 2024) introduces per-Gaussian embeddings and temporal embeddings. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Tian, Du, and Duan 2023; Gao et al. 2024), tracks (Wang et al. 2024; Seidenschwarz et al. 2024), and generative models (Wu et al. 2025; Chu, Ke, and Fragkiadaki 2024).
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works (Lee et al. 2023a; Ma et al. 2022; Wang et al. 2023; Lee et al. 2023b, 2024a; Zhao, Wang, and Liu 2024; Peng et al. 2024a; Lee et al. 2024b; Chen and Liu 2024; Oh et al. 2024) suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DyBluRF (Sun et al. 2024a; Bui et al. 2023) incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
@@
Motion blur occurs due to camera shake and object movement, which can be regarded as the integration of latent sharp images (Nah, Hyun Kim, and Mu Lee 2017), i.e.,
@@
$\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works (Zhao, Wang, and Liu 2024; Sun et al. 2024a; Wang et al. 2023) divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
@@
To estimate continuous camera poses, recent methods (Zhao, Wang, and Liu 2024; Peng et al. 2024a; Chen and Liu 2024; Oh et al. 2024) directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$),i.e.,
@@
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in the Sec.A of Suppl) for more stable optimization. We pre-train it and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
@@
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest varying the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps, thus reducing the difficulty of motion modeling. One way to achieve this is to uniformly divide the video into $L$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $L$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  (Bansal, Raj, and Choudhury 2016; Ren, Qian, and Chen 2020), i.e.,
@@
Second, despite different motions, the content of multiple frames within exposure time should be similar. We utilize $\mathcal{L}_{fc}$ to constrain consistency between neighbor frames, and that between each frame and the first frame, i.e.,
@@
Third, the blur in the lower resolution is lower level and is easier to remove (Kim, Lee, and Cho 2022; Tao et al. 2018), thus the artifacts are less in models trained with down-sampled blurry video. Taking this advantage, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models,i.e.,
@@
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion (Wang et al. 2024), we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details are in the Sec.B of Suppl.
@@
Table 1: novel view synthesis results on synthetic videos.
@@ Methods
Table 2: novel view synthesis results on real-world videos.
@@
Training Details. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $400$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ follows Shape-of-Motion (Wang et al. 2024). $N$ is set to 11. $L$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. All experiments are conducted with PyTorch (Paszke et al. 2019) on one Nvidia GeForce RTX A6000 GPU.
Evaluation Configurations. The synthetic data contains 9 scenes with significant motion blur from Stereo Blur Dataset (Zhou et al. 2019), where each scene contains blurry stereo videos and the corresponding sharp ones. Note that DyBluRF (Sun et al. 2024a) conducts experiments on $\times 2.5$ down-sampled data. We evaluate on both $\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$). For novel-view synthesis and deblurring, the rendering results may be spatially misaligned with ground truth due to the calibration error of camera parameters. Thus, we first freeze the pre-trained 4D model and optimize camera poses by minimizing the photometric error between rendering results and ground truth, and then calculate metrics (i.e., PSNR (Huynh-Thu and Ghanbari 2008), SSIM (Wang et al. 2004) and LPIPS (Zhang et al. 2018b)), following COLMAP-Free 3DGS (Fu et al. 2024). As there is no ground truth for frame interpolation and video stabilization, we employ recent no-reference metrics, i.e., CLIPIQA (Wang, Chan, and Loy 2023) and MUSIQ (Ke et al. 2021). Besides, we evaluate on 6 real-world blurry videos (i.e., Redmi data) captured by a Redmi K50 Ultra smartphone and 12 real-world ones (i.e., BARD-GS data) from BARD-GS (Lu et al. 2025), where each one contains 24 frames. For novel-view synthesis, we employ no-reference metrics (i.e., CLIPIQA and MUSIQ) and full-reference metrics (i.e., PSNR, SSIM and LPIPS) for the two data, respectively. For the other three tasks, we use no-reference metrics (i.e., CLIPIQA and MUSIQ) due to no ground truth.
@@
We compare with 7 state-of-the-art methods (i.e., DeformableGS (Yang et al. 2024), 4DGaussians (Wu et al. 2024), E-D3DGS (Bae et al. 2024), Shape-of-Motion (Wang et al. 2024), SplineGS (Park et al. 2025), DyBluRF (Sun et al. 2024a) and BARD-GS (Lu et al. 2025)), where DyBluRF and BARD-GS are designed for 4D reconstruction from blurry monocular video based on NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) respectively.
@@
Shape-of-Motion
None 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
@@
Novel-view synthesis. table 1 and table 2 summarize the results. First, methods (i.e., DyBluRF and BARD-GS) that perform 4D reconstruction and motion blur modeling jointly yield overall better performance, especially in LPIPS score. Although SplineGS (Park et al. 2025) gets better PSNR and SSIM scores in BARD-GS data, it produces blurry outputs. It is consistent with the finding in BARD-GS (Wang et al. 2024) that PSNR can sometimes yield higher values even when images appear blurrier. Second, benefiting the explicit 3D representation manner, BARD-GS outperforms DyBluRF, being the most competitive method. Third, compared with BARD-GS, our Deblur4DGS performs better due to the introduction of a series of regularization terms to avoid trivial solutions and blur-aware variable canonical Gaussians to better represent dynamic objects. Visual results in fig. 9 shows that Deblur4DGS removes blur more clearly and produces less visual artifacts in both static and dynamic areas. Per-scene results, more visual results and rendering speed comparisons are in the Sec.C of Suppl.
@@
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang, Xie, and Yao 2024)) deblurring methods and then perform 4D reconstruction, as shown in table 3. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results.
@@
Deblurring. Apart from 4D reconstruction-based methods, we compare with some state-of-the-art image and video deblurring ones. The results are in the Sec.C of Suppl. Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to deblurring-specific ones. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in supervised manner while Deblur4DGS is optimized with the given blurry video in a self-supervised manner. Although the data prior makes them perform better, Deblur4DGS is more convenient to use.
@@
Frame Interpolation. We interpolate camera poses and timestamps to generate $\times 16$ frame interpolation results. We compare with 4D reconstruction-based methods and some video frame interpolation ones (i.e., RIFE (Huang et al. 2022), EMAVFI (Zhang et al. 2023a), and VIDUE (Shang et al. 2023)). The results are in the Sec.C of Suppl. VIDUE is trained with large paired data in a supervised manner for joint deblurring and fame interpolation, thus achieving better results. Deform4DGS performs satisfactorily overall.
@@
Video Stabilization. We employ a Gaussian filter to smooth camera poses for video stabilization, following  (Peng et al. 2024b). The results are in the Sec.C of Suppl. Deblur4DGS achieves pleasant scores compared with 2D video stabilization methods (i.e., MeshFlow (Liu et al. 2016) and NNDVS (Zhang et al. 2023b)) and 4D reconstruction-based ones, which benefits from better geometry reconstruction.
@@
The results are summarized in table 6. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates this, leading to performance gain. We also experiment with an optical flow-based strategy (Shaw et al. ) to select canonical Gaussians, performs similar to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians corresponding to the sharper frame help blur removal. Visual results are in the Sec.D of Suppl.

### v3

@@ Abstract
Recent 4D reconstruction methods have yielded impressive results but rely on sharp videos as supervision. However, motion blur often occurs in videos due to camera shake and object movement, while existing methods render blurry results when using such videos for reconstructing 4D models. Although a few approaches attempted to address the problem, they struggled to produce high-quality results, due to the inaccuracy in estimating continuous dynamic representations within the exposure time. Encouraged by recent works in 3D motion trajectory modeling using 3D Gaussian Splatting (3DGS), we take 3DGS as the scene representation manner, and propose Deblur4DGS to reconstruct a high-quality 4D model from blurry monocular video. Specifically, we transform continuous dynamic representations estimation within an exposure time into the exposure time estimation. Moreover, we introduce the exposure regularization term, multi-frame, and multi-resolution consistency regularization term to avoid trivial solutions. Furthermore, to better represent objects with large motion, we suggest blur-aware variable canonical Gaussians. Beyond novel-view synthesis, Deblur4DGS can be applied to improve blurry video from multiple perspectives, including deblurring, frame interpolation, and video stabilization. Extensive experiments in both synthetic and real-world data on the above four tasks show that Deblur4DGS outperforms state-of-the-art 4D reconstruction methods. The codes are available at https://github.com/ZcsrenlongZ/Deblur4DGS.
@@ Introduction
Substantial efforts have been made for 4D reconstruction, which has extensive applications in augmented reality and virtual reality. To model static scenes, Neural Radiance Field (NeRF) (Mildenhall et al. 2021) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) propose implicit neural representation manner and explicit Gaussian ellipsoids one, respectively. To model dynamic objects, implicit neural fields (Zhu et al. 2024; Yang et al. 2024; Wu et al. 2024; Yan et al. 2023) and explicit deformation (Duan et al. 2024b; Chu et al. 2024; Katsumata et al. 2024; Lin et al. 2024; Li et al. 2024; Wang et al. 2024) are suggested for motion representation. While achieving great progress, most methods rely on synchronized multi-view videos. They yield unsatisfactory results when applied to monocular video, where dynamic objects are only observed once at each timestamp. To alleviate the under-constrained nature of the problem, recent studies have introduced data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Zhu et al. 2024), tracks (Seidenschwarz and et al. 2024; Lei et al. 2024), and generative models (Wu et al. 2025; Chu et al. 2024; Zeng et al. 2025) for better 4D reconstruction.
Unfortunately, motion blur often arises due to camera shake and object movement. When reconstructing the 4D scene from the blurry video, the above methods usually render blurry results. The first step to solving this problem is to deal with camera motion blur, which is relatively simple. Some NeRF-based (Lee et al. 2023a; Wang et al. 2023b; Lee et al. 2023b) and 3DGS-based (Zhao et al. 2024; Chen and Liu 2024; Oh et al. 2024) methods have suggested jointly optimizing 3D representation and camera poses within the exposure time by calculating the reconstruction loss between the synthetic blurry images and the input blurry frames. In contrast, the object motion blur is more challenging to address, as the solution has to estimate continuous and sharp dynamic representations within the exposure time to simulate blurry frames.
In this work, we take 3DGS (Kerbl et al. 2023) as the scene representation manner to explore the problem, driven by two main motivations. First, its successful application in 4D reconstruction make this method highly promising. Second, the explicit 3D motion modeling presents an opportunity to simplify the complex continuous dynamic representations estimation within the exposure time into exposure time estimation, avoiding complex extra motion modeling in DyBluRF (Sun et al. 2024a; Bui and et al. 2023). Once the exposure time is estimated, continuous dynamic representations can be obtained by directly interpolating between representations at the nearest integer timestamps. We note that the concurrent work BARD-GS (Lu et al. 2025) adopts a similar strategy, but they perform unsatisfactorily due to the under-constrained optimization, especially for large object motion.
@@
Deep learning-based image (Ren et al. 2023; Li and et al. 2023; Wang et al. 2022; Zhang et al. 2022, 2024) and video (Zhong et al. 2023b; Pan et al. 2023; Zhong et al. 2023a, 2020; Chan et al. 2022) deblurring methods have been widely explored. Compared to image deblurring methods, video ones leverage temporal clues between consecutive frames for more effective restoration. DSTNet (Pan et al. 2023) develops a deep discriminative spatial and temporal network. BasicVSR++ (Chan et al. 2022) improves feature fusion with second-order feature propagation and flow-guided alignment. BSSTNet (Zhang and et al. 2024) introduces a blur map to sufficiently utilize the entire video, achieving recent state-of-the-art. When reconstructing from a blurry video, pre-processing it with the 2D deblurring method is a straightforward manner. However, 2D deblurring methods cannot perceive 3D structures and maintain scene geometric consistency, leading to unsatisfactory scene reconstruction.
@@
To reconstruct 3D models, NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) introduce implicit neural representation manner and explicit Gaussian ellipsoids one respectively, where the latter generally achieves better results. To reconstruct 4D models, most works (Somraj et al. 2024; Duan et al. 2024a; Lu et al. 2024; Lin et al. 2024; Li et al. 2024; Sun et al. 2024b; Wu et al. 2024; Yang et al. 2024; Mihajlovic et al. 2024; Wang and et al. 2025) incorporate implicit neural fields and explicit deformation for motion representation. Moreover, to better reconstruct from monocular video, some studies enhance 4D reconstruction with data-driven priors, such as depth maps (Lee et al. 2023c; Yang et al. ), optical flows (Gao et al. 2024; Wang and et al. 2025), tracks (Wang et al. 2024; Seidenschwarz and et al. 2024), and generative models (Wu et al. 2025; Chu et al. 2024). For example, GFlow (Wang and et al. 2025) utilizes only 2D priors to lift a video to a 4D scene. GaussianMarbles (Stearns et al. 2024) reduces the degrees of freedom of each Gaussian.
Note that these methods heavily rely on high-quality sharp videos for supervision and perform poorly when facing blurry inputs. To process camera motion in static areas, recent works (Lee et al. 2023a; Ma et al. 2022; Wang et al. 2023b; Lee et al. 2023b, 2024a; Zhao et al. 2024; Peng et al. 2024a; Lee et al. 2024b; Chen and Liu 2024; Oh et al. 2024) suggest jointly optimizing the scene representation and recovering the camera poses within the exposure time. To process object motion blur in dynamic scenes, DyBluRF (Sun et al. 2024a; Bui and et al. 2023) incorporates object motion blur formation into dynamic model optimization but faces challenges in producing high-quality images and achieving real-time rendering. In this work, with 3DGS as the scene representation manner, we develop Deblur4DGS to reconstruct a high-quality 4D model from a blurry video.
@@
Motion blur occurs due to camera shake and object movement, which can be regarded as the integration of latent sharp images (Nah and et al. 2017), i.e.,
@@
$\mathbf{B}\in\mathbb{R}^{H\times W\times 3}$ is the blurry image and $\mathbf{I}_{t}$ is the latent sharp one at $t$-th timestamp. $(u,v)$ is pixel location, $\tau$ is the camera exposure time, and $\phi$ is a normalization factor. To approximate the integral operation, recent works (Zhao et al. 2024; Sun et al. 2024a; Wang et al. 2023b) divide the exposure time into $N$ timestamps and regard the blurry image as the average of $N$ sharp images, i.e.,
@@
To estimate continuous camera poses, recent methods (Zhao et al. 2024; Peng et al. 2024a; Chen and Liu 2024; Oh et al. 2024) directly optimize exposure start and end poses (i.e., $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$). Then, the linear interpolation is performed between $\mathbf{P}_{t,1}$ and $\mathbf{P}_{t,N}$ to obtain the camera pose at $i$-th intermediate timestamp (i.e., $\mathbf{P}_{t,i}$),i.e.,
@@
We follow the manner but deploy a tiny MLP as the camera motion predictor (see details in the Suppl) for more stable optimization. We pre-train it and static Gaussians $\mathbf{S}$ with static reconstruction loss $\mathcal{L}_{rec}^{s}$, i.e.,
@@
Blur-Aware Variable Canonical Gaussians. Existing 4D reconstruction methods generally select a single canonical Gaussians $\mathbf{C}$ across the entire video, which may produce results with missing details in scenes with large motion. To alleviate the issue, we suggest varying the canonical Gaussians as time progresses. In such case, the $k$-th canonical Gaussians $\mathbf{C}_{k}$ is only used for some nearby timestamps, thus reducing the difficulty of motion modeling. One way to achieve this is to uniformly divide the video into $K$ segments and select $\mathbf{C}_{k}$ for $k$-th segment. Although it improves performance, selecting the one corresponding to the sharper frame is better for blur removal. In particular, we first uniformly divide the video into $K$ segments and calculate the blur level $b_{t}$ of dynamic areas for $t$-th frame following  (Bansal et al. 2016; Ren et al. 2020), i.e.,
@@
Second, despite different motions, the content of multiple frames within exposure time should be similar. We utilize $\mathcal{L}_{mfc}$ to constrain consistency between neighbor frames, and that between each frame and the first frame, i.e.,
@@
Third, the blur in the lower resolution is lower level and is easier to remove (Kim et al. 2022; Tao et al. 2018), thus the artifacts are less in models trained with down-sampled blurry video. Taking this advantage, we impose $\mathcal{L}_{mrc}$ to assist the optimization of high-resolution models with results from low-resolution models,i.e.,
@@
$\lambda_{e}$, $\lambda_{mfc}$, and $\lambda_{mrc}$ are set to 0.1, 2, 1, respectively. Besides, following Shape-of-Motion (Wang et al. 2024), we also use some other regularization terms $\mathcal{L}_{oth}$ to help reconstruct 3D motion better, and the details are in the Suppl.
@@ Methods
Table 1: novel view synthesis results on synthetic videos.
@@
Table 2: novel view synthesis results on real-world videos.
@@
Training Details. For stable optimization, we pre-train the camera motion predictor and static Gaussians $\mathbf{S}$ for $400$ epochs. After that, we jointly optimize the camera motion predictor, $\mathbf{S}$, canonical dynamic Gaussians $\{\mathbf{C}_{k}\}_{k=1}^{K}$, deformable operation $\mathcal{F}$ and exposure time parameters $\{\mathbf{w}_{t}\}_{t=1}^{T}$ for $200$ epochs. The learning rate for camera motion predictor is set to $5\times 10^{-4}$ and decayed to $1\times 10^{-5}$. The learning rate for $\{\mathbf{w}_{t}\}_{t=1}^{T}$ is set to $1\times 10^{-1}$ and decayed to $1\times 10^{-5}$. The learning rate for $\mathbf{S}$, $\{\mathbf{C}_{k}\}_{k=1}^{K}$ and $\mathcal{F}$ follows Shape-of-Motion (Wang et al. 2024). $N$ is set to 11. $K$ and $H$ are set to 5 and 3 respectively. $\mathbf{\epsilon}$ is set to $1.0$. Experiments are conducted with PyTorch (Paszke et al. 2019) on one Nvidia GeForce RTX A6000 GPU.
Evaluation Configurations. The synthetic data contains 9 scenes with significant motion blur from Stereo Blur Dataset (Zhou et al. 2019), where each scene contains blurry stereo videos and the corresponding sharp ones. Note that DyBluRF (Sun et al. 2024a) conducts experiments on $\times 2.5$ down-sampled data. We evaluate on both $\times 2.5$ down-sampled ones (i.e., $288\times 512$) and the original ones (i.e., $720\times 1080$). For novel-view synthesis and deblurring, the rendering results may be spatially misaligned with ground truth due to the calibration error of camera parameters. Thus, we first freeze the pre-trained 4D model and optimize camera poses by minimizing the photometric error between rendering results and ground truth, and then calculate metrics (i.e., PSNR, SSIM and LPIPS (Zhang et al. 2018)), following COLMAP-Free 3DGS (Fu et al. 2024). As there is no ground truth for frame interpolation and video stabilization, we employ recent no-reference metrics, i.e., CLIPIQA (Wang et al. 2023a) and MUSIQ (Ke and et al. 2021). Besides, we evaluate on 6 real-world blurry videos (i.e., Redmi data) captured by a Redmi K50 Ultra smartphone and 12 real-world ones (i.e., BARD-GS data) from BARD-GS (Lu et al. 2025), where each one contains 24 frames. For novel-view synthesis, we employ no-reference metrics (i.e., CLIPIQA and MUSIQ) and full-reference metrics (i.e., PSNR, SSIM and LPIPS) for the two data, respectively. For the other three tasks, we use no-reference metrics due to no ground truth.
@@
We compare with 7 state-of-the-art methods (i.e., DeformableGS (Yang et al. 2024), 4DGaussians (Wu et al. 2024), E-D3DGS (Bae et al. 2024), Shape-of-Motion (Wang et al. 2024), SplineGS (Park and et al. 2025), DyBluRF (Sun et al. 2024a) and BARD-GS (Lu et al. 2025)), where DyBluRF and BARD-GS are designed for 4D reconstruction from blurry monocular video based on NeRF and 3DGS respectively.
@@
Shape-of-Motion None 26.06 / 0.910 / 0.144 25.81 / 0.897 / 0.246
@@
Novel-view synthesis. table 1 and table 2 summarize the results. First, methods (i.e., DyBluRF and BARD-GS) that perform 4D reconstruction and motion blur modeling jointly yield overall better performance, especially in LPIPS score. Although SplineGS gets better PSNR and SSIM scores in BARD-GS data, it produces blurry outputs. It is consistent with the finding in BARD-GS (Wang et al. 2024) that PSNR can sometimes yield higher values even when images appear blurrier. Second, benefiting the explicit 3D representation manner, BARD-GS outperforms DyBluRF, being the most competitive method. Third, compared with BARD-GS, our Deblur4DGS performs better due to the introduction of a series of regularization terms to avoid trivial solutions and blur-aware variable canonical Gaussians to better represent dynamic objects. Visual results in fig. 11 and fig. 2 shows that Deblur4DGS removes blur more clearly and produces less visual artifacts in both static and dynamic areas. Per-scene results and more visual results are in the Suppl.
@@
In addition, to further demonstrate the effectiveness of Deblur4DGS, we first pre-process the blurry videos with state-of-the-art image (i.e., Restormer (Zamir et al. 2022)) or video (DSTNet (Pan et al. 2023) and BSSTNet (Zhang and et al. 2024)) deblurring methods and then perform 4D reconstruction, as shown in table 3. Compared with reconstruction from blurry videos, the incorporation of deblurring models improves performance. This is because the deblurring models remove some blur, facilitating sharp scene reconstruction. However, as the deblurring methods cannot perceive 3D structure and maintain scene geometric consistency, the reconstruction results are still unsatisfactory. In contrast, Deblur4DGS jointly reconstructs scene geometry and processes motion blur in 3D space, achieving better scene reconstruction results. Visual results are in the Suppl.
@@
Deblurring. Apart from 4D reconstruction-based methods, we compare with some state-of-the-art image and video deblurring ones. The results are in the Suppl. Deblur4DGS obtains better results than 4D reconstruction-based methods and comparable ones to deblurring-specific ones. Compared with the former, Deblur4DGS better reconstructs the scene, thus performing better. Note that the latter ones are trained on large paired data in supervised manner while Deblur4DGS is optimized with the given blurry video in a self-supervised manner. Although the data prior makes them perform better, Deblur4DGS is more convenient to use.
@@
Frame Interpolation. We interpolate camera poses and timestamps to generate $\times 16$ frame interpolation results. We compare with 4D reconstruction-based methods and some video frame interpolation ones (i.e., RIFE (Huang et al. 2022), EMAVFI (Zhang et al. 2023a), and VIDUE (Shang et al. 2023)). The results are in the Suppl. VIDUE is trained with large paired data in a supervised manner for joint deblurring and fame interpolation, thus achieving better results.
@@
Video Stabilization. We employ a Gaussian filter to smooth camera poses for video stabilization, following  (Peng et al. 2024b). The results are in the Suppl. Deblur4DGS achieves pleasant scores compared with 2D video stabilization methods (i.e., MeshFlow (Liu et al. 2016) and NNDVS (Zhang et al. 2023b)) and 4D reconstruction-based ones, which benefits from better geometry reconstruction.
@@
The results are summarized in table 6. First, selecting a single canonical Gaussians across the entire video (i.e., None) leads to poor performance, due to the challenge of modeling large object motion. Second, selecting variable canonical Gaussians uniformly (i.e., w/o Blur-Aware) alleviates this, leading to performance gain. We also experiment with an optical flow-based strategy (Shaw et al. ) to select canonical Gaussians, performs similar to the uniform selection. This may be due to the inaccurate estimation of optical flow from blurry images. Third, our blur-aware selection is better, as the canonical Gaussians from the sharper frame help blur removal. Visual results are in the Suppl.